OpenStudy (anonymous):

Please, help me to solve this: abs(x^2-1)>=abs(2x)-1/3 (algebraically ) sorry

OpenStudy (anonymous):

\[\left| x^2-1 \right|\ge \left| 2x \right|-\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

http://fooplot.com/plot/bapz7l4kw9

OpenStudy (phi):

based on the graph, there are 3 regions where the red line is above the green line those are your answer

OpenStudy (anonymous):

algebraically ?

OpenStudy (phi):

we would have to break the problem into pieces. notice that x^2 - 1 is always positive when x^2 ≥ 1, and so is 2x so for x≥1 we can solve x^2 -1 ≥ 2x - 1/3 or x^2 -2x - 2/3 ≥ 0 (x - 1 + sqrt(5/3))( x -1 - sqrt(5/3)) ≥ 0 x≥ 1 + sqrt(5/3) is one answer. Based on the graph, I think we will find -1 - sqrt(5/3) will be another, but we can check by assuming x ≤ -1 and solving x^2 -1 ≥ -2x - 1/3 also, check the region between 0 and +1 and the region between -1 and 0

OpenStudy (anonymous):

phi, I'll try to study your answer. Thank you for the attention.